OpenStudy (anonymous):

HELP ME!!! Explain why the initial value of any function of the form f(x) = a(bx) is equal to a.

OpenStudy (er.mohd.amir):

if x start from x=0,1,2,3......... then f(0)=0 f(1)=ab and so on so f(x) depends on value of x .

OpenStudy (er.mohd.amir):

or f(x)=a+bx then initial value is a due to f(0)=a+b*0 =a+0 =a

OpenStudy (er.mohd.amir):

U write Q is wrong.